find a general linear equation of the straight line that passes through (-5/2, 5) and has a slope of 1/3?
Answer by mananth(16946)

m= 1/3
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( 2.5 , 5 ) in
Y = m x + b
5.00 = 1/3 * 5/2 + b
5.00 = 5/6 + b
b= 5 + - 5/6
b= 25/ 6
So the equation will be
Y = 1/3 x + 25/6
m.ananth@hotmail.ca